Two protons are separated by a distance of 3.6×10^10 m in air

A. find the magnitude of the electric force once protons exerts on other. is the force attractive or repulsive?

B. find the magnitude of the gravitational force once proton is exerts on the other. is this force attractive or repulsive?



1
Expert's answer
2022-03-02T14:39:16-0500

a) The electric force (repulsive, since protons are both positively charged):


"F = k\dfrac{e^2}{r^2} = 9\times 10^9\cdot\dfrac{(1.6\times 10^{-19})^2}{(3.6\times 10^{10})^2}\approx 1.8\times 10^{-49}N"

b) The gravitational force (always attractive):


"F = G\dfrac{m^2}{r^2} = 6.67\times 10^{-11}\cdot \dfrac{(1.67\times 10^{-27})^2}{(3.6\times 10^{10})^2}\approx 1.4\times 10^{-85}N"